useing a mid-point formula Line AB is located (-5,0) and (-1,3).what slope is
a) 4/3
b) 2/4
c) 3/4
d) 4/2

Answer by checkley77(12844)

slope=(3-0)/(-1+5)=3/4
Answer:C)=3/4